Seven years ago, Tom was 5 times as old as Betty was. Four years ago, Tom was only 3 times as old as Betty was. Find the present age of each person.
I really need help to set up the equation. I tried ;
T-7=5(b-7)
T-4=3(b-4)
But when I went to solve the answer was totally off. Please help asap!

Answer by Alan3354(69443)

Seven years ago, Tom was 5 times as old as Betty was. Four years ago, Tom was only 3 times as old as Betty was. Find the present age of each person.
I really need help to set up the equation. I tried ;
T-7=5(b-7)
T-4=3(b-4)
====================
Your equations are good.
T-7=5(b-7)
T-4=3(b-4)
-----
t-7 = 5b-35
t = 5b - 28
----
t-4 = 3b-12
t = 3b - 8
==================
t = 5b - 28
t = 3b - 8
Since both RH sides = t
5b - 28 = 3b - 8
2b = 20
b = 10
---
t = 22